OpenStudy (anonymous):

Find the area of the circle in simplest polynomial form. A = 3.14r2 2x+1 a: 6.28x2 + 3.14 b: 12.56x2 + 3.14 c: 6.28x2 + 3.14x + 3.14 d: 12.56x2 + 12.56x + 3.14

OpenStudy (anonymous):

What one you think?

OpenStudy (anonymous):

A

OpenStudy (nurali):

A = 3.14r^2 since r = 2x+1 A = 3.14(2x+1)^2 Now all you have to do is simplify this expression. A=3.14(2x+1)(2x+1) A=3.14(4x^2 + 2x + 2x + 1) A=3.14(4x^2 +4x+1) A=12.56x^2 + 12.56x + 3.14

OpenStudy (welshfella):

Oh yes - i didn't understand the question at first you didnt say that r = 2x + 1

OpenStudy (anonymous):

i don't know how to do that

OpenStudy (welshfella):

the answers not A

OpenStudy (welshfella):

Nurali has worked it out for you

OpenStudy (welshfella):

you just say A = 3.14 r^2 when r = 2x + 1

OpenStudy (anonymous):

is it B

OpenStudy (anonymous):

i really don't understand how to do this stuff

OpenStudy (anonymous):

hello?

OpenStudy (welshfella):

its basic algebra you replace the r in 3.14 r^2 by 2x + 1 and simplify